How can i find depth of an object in water when i'm given the density and pressure?
1
Expert's answer
2019-11-26T10:35:47-0500


We use the formula for the pressure of the column of fluid height "h"

"p=\rho \cdot g \cdot h"

Where "\rho -" fluid density "p -" deep pressure "h"

"g-" acceleration of gravity

then write

"h=\frac{p}{\rho \cdot g}"